The efficiency of a wheel and axle system is 80% and the ratio of wheel to radius of axle is 4:1. In order to lift a mass of 20k, the effort required is?


1
Expert's answer
2021-06-01T17:38:44-0400

"\eta \frac Rr=\frac NF, \implies"

"F=\frac{N}{\eta \frac Rr}=\frac{mg}{\eta \frac Rr},"

"F=\frac{20\cdot 10}{0.8\cdot (4:1)}=62.5~\text N."